OpenStudy (keke):

simplify; 2 over √(5)

OpenStudy (anonymous):

\[\frac{2}{\sqrt{5}}=\frac{2\sqrt{5}}{5}\]

OpenStudy (anonymous):

well u can multiply top and bottom by sqrt5

OpenStudy (anonymous):

as bendt as done